# A Set of Tutorials for the LAMMPS Simulation Package [Article v1.0]

### v2

@@ Abstract
The availability of open-source molecular simulation software packages allows scientists and engineers to focus on running and analyzing simulations without having to write, parallelize, and validate their own simulation software. While molecular simulations thus become accessible to a larger audience, the “black box” nature of such software packages and wide array of options and features can make it challenging to use them correctly, particularly for beginners in the topic of MD simulations. LAMMPS is one such versatile molecular simulation code, designed for modeling particle-based systems across a broad range of materials science and computational chemistry applications, including atomistic, coarse-grained, mesoscale, grid-free continuum, and discrete element models. LAMMPS is capable of efficiently running simulations of varying sizes from small desktop computers to large-scale supercomputing environments. Its flexibility and extensibility make it ideal for complex and extensive simulations of atomic and molecular systems, and beyond. This article introduces a suite of tutorials designed to make learning LAMMPS more accessible to new users. The first four tutorials cover the basics of running molecular simulations in LAMMPS with systems of varying complexities. The second four tutorials address more advanced molecular simulation techniques, specifically the use of a reactive force field, grand canonical Monte Carlo, enhanced sampling, and the REACTER protocol. In addition, we introduce LAMMPS–GUI, an enhanced cross-platform graphical text editor specifically designed for use with LAMMPS and able to run LAMMPS directly on the edited input. LAMMPS–GUI is used as the primary tool in the tutorials to edit inputs, run LAMMPS, extract data, and visualize the simulated systems.
@@ 1.1 Scope
In Tutorial 5, the ReaxFF reactive force field is used, specifically designed to simulate chemical reactions by dynamically adjusting atomic interactions [14]. ReaxFF includes charge equilibration (QEq), a method that allows the partial charges of atoms to adjust according to their local environment.
@@ 3.1.1 My first input
The first line, region simbox (…), defines a region named simbox that is a block (i.e., a rectangular cuboid) extending from -20 to 20 units along all three spatial dimensions. The second line, create_box 2 simbox, initializes a simulation box based on the region simbox and reserves space for two types of atoms. In LAMMPS, every atom is assigned an atom type property. This property selects which force field parameters (here, the Lennard-Jones parameters, and ) are applied to each pair of atoms via the pair_coeff command. We discuss in Tutorial 2 how this applies to many-body pair styles, and in Tutorial 3 how this applies to Coulomb interactions.
@@
By default, LAMMPS calculates the mixed force field coefficients for different atom types using geometric averages: , . In the present case, , and . Other rules can be selected using the pair_modify command.
@@
You can now run LAMMPS (see subsection A.3 for details on running LAMMPS). The simulation should finish quickly, and with the default settings, LAMMPS–GUI will open two windows: one displaying the console output and another with a chart. The «Output» window will display information from the executed commands, including the total energy and pressure at step 0, as specified by the thermodynamic data request. Since no actual simulation steps were performed, the «Charts» window will be empty.
@@
The potential energy, , decreases from a positive value to a negative value (Figs. 4 and 5 a). Note that during energy minimization, the potential energy equals the total energy of the system, , since the kinetic energy, , is zero. The initially positive potential energy is expected, as the atoms are created at random positions within the simulation box, with some in very close proximity to each other. This proximity results in a large initial potential energy due to the repulsive branch of the Lennard-Jones potential [i.e., the term in Eq. (1)]. As the energy minimization progresses, the energy decreases - first rapidly - then more gradually, before plateauing at a negative value. This indicates that the atoms have moved to reasonable distances from one another.
@@
After energy minimization, any overlapping atoms are displaced, and the system is ready for a molecular dynamics simulation. To continue from the result of the minimization step, append the MD simulation commands to the same input script, initial.lmp. Add the following lines immediately after the minimize command:
@@ 3.1.2 Improving the script
The side in and side out keywords are used to define regions representing the inside and outside of the cylinder of radius 10 length units. Then, append a sixth section titled Save system at the end of the file, ensuring that the write_data command is placed after the minimize command:
@@
LAMMPS compute commands can produce a wide variety of data and one can identify the category from the name of the compute style: global data (no suffix), local data (/local suffix), per-atom data (/atom suffix), per-chunk data (/chunk suffix), per-gridpoint data (/grid suffix). In the example above, the compute coord/atom produces per-atom data, which is used as input for compute reduce which returns global data. For global data three kinds of data exists: scalars (single values), vectors (one-dimensional arrays), or arrays (two-dimensional tables). When referencing results of a compute, you can use indices: for example, c_mycompute refers to the entire scalar, vector, or array, and c_mycompute[1] refers to its first element (in case of vector or array). In some cases also wildcards like “*” can be used to, for instance, refer to all elements of a vector instead of having specify all elements individually. In general, “consumer” commands (fix styles or dump styles, variables, or other compute styles) can only work with certain data types or need to have keywords set to select which data to use. You need to check the documentation of each command to ensure compatibility.
@@ 3.2.1 Unbreakable bonds
The pair_coeff command sets the parameters for non-bonded Lennard-Jones interactions atom type 1 to and . The bond_coeff provides the equilibrium distance and the spring constant for the harmonic potential imposed between two neighboring carbon atoms. The potential is given by . The angle_coeff gives the equilibrium angle and constant for the potential between three neighboring atoms : . The dihedral_coeff and improper_coeff define the potentials for the constraints between 4 atoms.
@@
With the thermo_modify command, we specify to LAMMPS that the temperature of the middle group, cnt_mid, must be outputted, instead of the temperature of the entire system. This choice is motivated by the presence of frozen parts with an effective temperature of 0 K, which makes the average temperature of the entire system less relevant. The thermo_modify command also imposes the use of the YAML format that can easily be read by Python (see below).
@@ 3.3 Tutorial 3: Polymer in water
The goal of this tutorial is to use LAMMPS to solvate a small hydrophilic polymer (PEG - polyethylene glycol) in a reservoir of water (Fig. 16). Once the water reservoir is properly equilibrated at the desired temperature and pressure, the polymer molecule is added and a constant stretching force is applied to both ends of the polymer. The evolution of the polymer length is measured as a function of time. The GROMOS 54A7 force field [36] is used for the PEG, the SPC/Fw model [37] is used for the water, and the long-range Coulomb interactions are solved using the PPPM solver [38]. This tutorial was inspired by a publication by Liese and coworkers, in which molecular dynamics simulations are compared with force spectroscopy experiments, see Ref. 39.
@@ 3.3.1 Preparing the water reservoir
The first parameter is 0, meaning that the atom IDs from the water.mol file will be used. The overlap 1.0 option of the create_atoms command ensures that no atoms are placed exactly in the same position, as this would cause the simulation to crash. The maxtry 50 asks LAMMPS to try at most 50 times to insert the molecules, which is useful in case some insertion attempts are rejected due to overlap. In some cases, depending on the system and the values of overlap and maxtry, LAMMPS may not create the desired number of molecules. Always check the number of created atoms in the log file (or in the «Output» window), where you should see:
@@ 3.4.1 System preparation
Within the last three lines, a region named rliquid is created based on the last defined lattice, fcc 4.04. rliquid will be used for introducing the water molecules. The molecule command opens up the molecule template called water.mol, and names the associated molecule h2omol. The new molecules are placed on the fcc 4.04 lattice by the create_atoms command. The first parameter is 0, meaning that the atom IDs from the water.mol file will be used. The number 482793 is a seed that is required by LAMMPS, it can be any positive integer.
@@
The run 0 command initializes the simulation but does not advance positions or velocities which is required for cleanly saving the state. The write_data command generates a file called system.data containing the information required to restart the simulation from the final configuration produced by this input file. With the nocoeff option, the parameters from the force field are not included in the .data file. Run the create.lmp file using LAMMPS, and a file named create.data will be created alongside create.lmp.
@@
The first two variables extract the centers of mass of the two walls. The deltaz variable is then used to calculate the difference between the two variables walltopz and wallbotz, i.e. the distance between the two centers of mass of the walls.
@@ 3.4.2 Imposed shearing
The setforce commands cancel the forces on walltop and wallbot in the direction. As a result, the atoms in these two groups will not experience any forces along from the rest of the system. Consequently, in the absence of external forces, these atoms will conserve the initial velocities imposed by the two velocity commands. As seen previously, although the forces on these atoms are set to zero, the fix setforce still stores the forces acting on the group before cancellation, which can later be extracted for analysis (see below).
@@ 3.6.1 Generation of the silica block
In line with what is done in previous tutorials, the create_atoms commands are used to place 240 Si atoms and 480 O atoms, respectively. This corresponds to an initial density of approximately g/cm3, which is close to the expected final density of amorphous silica at 300 K.
@@
Finally, let us implement the annealing procedure which consists of three consecutive runs. This procedure was inspired by Ref. [56]. First, to melt the system, a phase at is performed:
Next, a second phase, during which the system is cooled down from to , is implemented as follows:
n this case, the initial and final target temperatures set for the Nosé-Hoover thermostat is different, causing it to evolve linearly within the number of timesteps evoked in the run command. In the third step, the system is equilibrated at the final desired conditions, and , using an anisotropic pressure coupling:
@@ 3.6.2 Cracking the silica
As discussed, the fix nvt command integrates the Nosé-Hoover equations of motion to sample the NVT ensemble, which allows controlling the temperature of the system. As observed from the generated images, the atoms progressively adjust to the changing box dimensions. At some point, bonds begin to break, leading to the appearance of dislocations (Fig. 37).
@@ 3.6.3 Adding water
Here, a variable of type ‘atom’ is used. Such variable defines a per-atom property, i.e., it evaluates the specified expression separately for each atom. This is often used to select atoms based on their properties or types.
@@ 3.7.1 Method 1: Free sampling
where is the free energy difference, is the gas constant, is the temperature, is the density, and is a reference density. As an illustration, let us apply this method to a simple configuration that consists of a particles in a box in the presence of a position-dependent repulsive force that makes the center of the box a less favorable area to explore.
@@
Here, we begin by defining variables for the Lennard-Jones interaction and and for the repulsive potential , which are , , and [see Eqs. (4-5) below]. The cut-off value of was chosen to create a Weeks-Chandler-Andersen (WCA) potential, which is a truncated and purely repulsive LJ potential [64]. The potential is also shifted to be equal to 0 at the cut-off using the pair_modify command. The unit system is real, in which energy is in kcal/mol, distance in Ångströms, or time in femtosecond, has been chosen for practical reasons: the WHAM algorithm used in the second part of the tutorial automatically assumes the energy to be in kcal/mol.
@@
The variables , , and , defined in the previous subsection, are used here to create the repulsive potential, restricting the atoms from exploring the center of the box:
@@
Fig. 41 shows the potential and force along the -axis. With , is of the same order of magnitude as the thermal energy , where is the Boltzmann constant and is the temperature used in this simulation. Under these conditions, particles are expected to frequently overcome the energy barrier due to thermal agitation.
@@
Here, the chunk/atom command discretizes the simulation domain into spatial bins of size 2 Å along the direction, and the ave/chunk command outputs the number density of atoms within each bin to the file free-sampling.dat. The step count is reset to 0 using reset_timestep to synchronize it with the output times of fix density/number. Run the simulation using LAMMPS.
@@
Increasing the value of reduces the average number of atoms in the central region, making it difficult to achieve a high-resolution free energy profile within reasonable simulation times. For example, running the same simulation with , corresponding to , results in no atoms exploring the central part of the simulation box during the simulation. In such a case, employing an enhanced sampling method is recommended, as done in the next section.
@@ 3.7.2 Method 2: Umbrella sampling
The first difference from the previous case is the larger value for the repulsive potential , which makes the central area of the system very unlikely to be visited by free particles. The second difference is the introduction of the variable , which will be used for the biasing potential.
@@ 3.8.1 Creating the system
The class2 styles compute a 6/9 Lennard-Jones potential [70]. The class2 bond, angle, dihedral, and improper styles are used as well, see the documentation for a description of the respective potential form they, each, prescribe. The tail yes option adds long-range van der Waals tail corrections to the energy and pressure. The mix sixthpower imposes the following mixing rule for the calculation of the cross coefficients:
@@
These commands should be familiar from previous tutorials.
@@ 3.8.3 Simulating the reaction
Here, the thermo custom command is used to print the cumulative reaction counts which are calculated by fix rxn and thus can be extracted from it. Run the simulation using LAMMPS. As the simulation progresses, polymer chains are observed forming (Fig. 49). During this reaction process, the temperature of the system remains well-controlled (Fig. 50 a), while the number of reactions, , increases with time (Fig. 50 b).

### v3

@@ Abstract
The availability of open-source molecular simulation software packages allows scientists and engineers to focus on running and analyzing simulations without having to write, parallelize, and validate their own simulation software. While molecular simulations thus become accessible to a larger audience, the “black box” nature of such software packages and wide array of options and features can make it challenging to use them correctly, particularly for beginners in the topic of simulations. LAMMPS is one such versatile molecular simulation code, designed for modeling particle-based systems across a broad range of materials science and computational chemistry applications, including atomistic, coarse-grained, mesoscale, grid-free continuum, and discrete element models. LAMMPS is capable of efficiently running simulations of varying sizes from small desktop computers to large-scale supercomputing environments. Its flexibility and extensibility make it ideal for complex and extensive simulations of atomic and molecular systems, and beyond. This article introduces a suite of tutorials designed to make learning LAMMPS more accessible to new users. The first four tutorials cover the basics of running molecular simulations in LAMMPS with systems of varying complexities. The second four tutorials address more advanced molecular simulation techniques, specifically the use of a reactive force field, grand canonical Monte Carlo, enhanced sampling, and the REACTER protocol. In addition, we introduce LAMMPS–GUI, an enhanced cross-platform graphical text editor specifically designed for use with LAMMPS and able to run LAMMPS directly on the edited input. LAMMPS–GUI is used as the primary tool in the tutorials to edit inputs, run LAMMPS, extract data, and visualize the simulated systems.
@@ 1.1 Scope
In Tutorial 5, the ReaxFF reactive force field, which is specifically designed to simulate chemical reactions by dynamically adjusting atomic interactions [14], is used. ReaxFF includes charge equilibration (QEq), a method that allows the partial charges of atoms to adjust according to their local environment.
@@ 3.1.1 My first input
The first line, region simbox (…), defines a region named simbox that is a block (i.e., a rectangular cuboid) extending from -20 to 20 units along all three spatial dimensions. The second line, create_box 2 simbox, initializes a simulation box based on the region simbox and reserves space for two types of atoms. In LAMMPS, every atom is assigned an atom type property. This property selects which force field parameters (here, the Lennard-Jones parameters, and ) are applied to each pair of atoms via the pair_coeff command (see below). We discuss in Tutorial 2 how this applies to many-body pair styles, and in Tutorial 3 how this applies to Coulomb interactions.
@@
By default, LAMMPS calculates the mixed Lennard-Jones coefficients for pairs of atoms having distinct atom types using geometric averages: , . In the present case, , and . Other rules can be selected using the pair_modify command.
@@
You can now run LAMMPS (see subsection A.3 for details on running LAMMPS). The simulation should finish quickly, and, with the default settings, LAMMPS–GUI will open two windows: one displaying the console output and another with a chart. The «Output» window will display information from the executed commands, including the total energy and pressure at step 0, as specified by the thermodynamic data request. Since no actual simulation steps were performed, the «Charts» window will be empty.
@@
The potential energy, , decreases from a positive value to a negative value (Figs. 4 and 5 a). Note that during the energy minimization, the potential energy equals the total energy of the system, , since the kinetic energy, , is zero. The initially positive potential energy is expected, as the atoms are created at random positions within the simulation box, with some in very close proximity to each other. This proximity results in a large initial potential energy due to the repulsive branch of the Lennard-Jones potential [i.e., the term in Eq. (1)]. As the energy minimization progresses, the energy decreases - first rapidly - then more gradually, before plateauing at a negative value. This indicates that the atoms have moved to reasonable distances from one another.
@@
After the energy minimization, any overlapping atoms are displaced, and the system is ready for a molecular dynamics simulation. To continue from the result of the minimization step, append the MD simulation commands to the same input script, initial.lmp. Add the following lines immediately after the minimize command:
@@ 3.1.2 Improving the script
The side in and side out keywords are used to define regions representing the inside and outside of the cylinder of radius 10 length units, respectively. Then, append a sixth section titled Save system at the end of the file, ensuring that the write_data command is placed after the minimize command:
@@
LAMMPS compute commands can produce a wide variety of data and one can identify the category from the name of the compute style: global data (no suffix), local data (/local suffix), per-atom data (/atom suffix), per-chunk data (/chunk suffix), per-grid data (/grid suffix). In the example above, the compute coord/atom produces per-atom data, which is used as input for compute reduce which returns global data. For global data three kinds of data exists: scalars (single values), vectors (one-dimensional arrays), or arrays (two-dimensional tables). When referencing results of a compute, you can use indices: for example, c_mycompute refers to the entire scalar, vector, or array, and c_mycompute[1] refers to its first element or column (in case of vector or array). In some cases also wildcards like “*” can be used to, for instance, refer to all elements of a vector instead of having specify all elements individually. In general, “consumer” commands (fix styles or dump styles, variables, or other compute styles) can only work with certain data types or need to have keywords set to select which data to use. You need to check the documentation of each command to ensure compatibility.
@@ 3.2.1 Unbreakable bonds
The pair_coeff command sets the parameters for non-bonded Lennard-Jones interactions between atoms type 1 to and . The bond_coeff provides the equilibrium distance and the spring constant for the harmonic potential imposed between two bonded carbon atoms. The potential is given by . The angle_coeff gives the equilibrium angle and constant for the potential between atoms forming an angle: . The dihedral_coeff and improper_coeff define the potentials for the constraints between 4 atoms.
@@
With the thermo_modify command, we specify to LAMMPS that the temperature mid of the middle group, cnt_mid, must be outputted, instead of the temperature of the entire system. This choice is motivated by the presence of frozen parts with an effective temperature of 0 K, which makes the average temperature of the entire system less relevant. The thermo_modify command also imposes the use of the YAML format that can easily be read by Python (see below).
@@ 3.3 Tutorial 3: Polymer in water
The goal of this tutorial is to use LAMMPS to solvate a small hydrophilic polymer molecule (PEG - polyethylene glycol) in a reservoir of water (Fig. 16). Once the water reservoir is properly equilibrated at the desired temperature and pressure, the polymer molecule is added and a constant stretching force is applied to both ends of the polymer. The evolution of the polymer length is measured as a function of time. The GROMOS 54A7 force field [36] is used for the PEG, the SPC/Fw model [37] is used for the water, and the long-range Coulomb interactions are solved using the PPPM solver [38]. This tutorial was inspired by a publication by Liese and coworkers, in which molecular dynamics simulations are compared with force spectroscopy experiments, see Ref. 39.
@@ 3.3.1 Preparing the water reservoir
The first parameter is 0, meaning that the atom types from the water.mol file will be used. The overlap 1.0 option of the create_atoms command ensures that no atoms are placed exactly in the same position, as this would cause the simulation to crash. The maxtry 50 asks LAMMPS to try at most 50 times to insert the molecules, which is useful in case some insertion attempts are rejected due to overlap. In some cases, depending on the system and the values of overlap and maxtry, LAMMPS may not create the desired number of molecules. Always check the number of created atoms in the log file (or in the «Output» window), where you should see:
@@ 3.4.1 System preparation
Within the last three lines, a region named rliquid is created based on the last defined lattice, fcc 4.04. rliquid will be used for introducing the water molecules in the simulation domain. The molecule command opens up the molecule template called water.mol, and names the associated molecule h2omol. The new molecules are placed on the fcc 4.04 lattice by the create_atoms command. The first parameter is 0, meaning that the atom types from the water.mol file will be used. The number 482793 is a seed that is required by LAMMPS, it can be any positive integer.
@@
The run 0 command initializes the simulation, which is required for cleanly saving the state, but it does not advance positions or velocities. The write_data command generates a file called system.data containing the information required to restart the simulation from the final configuration produced by this input file. With the nocoeff option, the parameters from the force field are not included in the .data file. Run the create.lmp file using LAMMPS, and a file named create.data will be created alongside create.lmp.
@@
The first two variables extract the z coordinate of the centers of mass of the two walls. The deltaz variable is then used to calculate the difference between the two variables walltopz and wallbotz, i.e. the distance in the z direction between the two centers of mass of the walls.
@@ 3.4.2 Imposed shearing
The setforce commands cancel the forces on walltop and wallbot in the direction. As a result, the atoms in these two groups will not experience any forces along from their interaction with rest of the system. Consequently, in the absence of external forces, these atoms will conserve the initial velocities imposed by the two velocity commands. As seen previously, although the forces on these atoms are set to zero, the fix setforce still stores the forces acting on the group before cancellation, which can later be extracted for analysis (see below).
@@ 3.6.1 Generation of the silica block
In line with what is done in previous tutorials, the create_atoms commands are used to place 240 Si atoms and 480 O atoms, respectively, in the region previously defined. This corresponds to an initial density of approximately g/cm3, which is close to the expected final density of amorphous silica at 300 K.
@@
Finally, let us implement the annealing procedure which consists of three consecutive runs. This procedure was inspired by Ref. [56]. First, to melt the system, a run at is performed:
Next, a second run, during which the system is cooled down from to , is implemented as follows:
In this case, the initial and final target temperatures set for the Nosé-Hoover thermostat is different, causing it to evolve linearly within the number of timesteps evoked in the run command. In the third run, the system is equilibrated at the final desired conditions, and , using an anisotropic pressure coupling:
@@ 3.6.2 Cracking the silica
As discussed, the fix nvt command integrates the Nosé-Hoover equations of motion and is employed to control the temperature of the system. As observed from the generated images, the atoms progressively adjust to the changing box dimensions. At some point, bonds begin to break, leading to the appearance of dislocations (Fig. 37).
@@ 3.6.3 Adding water
Here, a variable of style ‘atom’ is used. Such variable defines a per-atom property, i.e., it evaluates the specified expression separately for each atom. This is often used to select atoms based on their properties or types.
@@ 3.7.1 Method 1: Free sampling
where is the free energy difference, is the gas constant, is the temperature, is the density, and is a reference density. As an illustration, let us apply this method to a simple system that consists of a particles in a box in the presence of a position-dependent repulsive force that makes the center of the box a less favorable area to explore.
@@
Here, we begin by defining variables for the Lennard-Jones parameters ( and ) and for the repulsive potential parameters , which are 0, , and [see Eqs. (4-5) below]. The cut-off value of was chosen to create a Weeks-Chandler-Andersen (WCA) potential, which is a truncated and purely repulsive LJ potential [64]. The potential is also shifted to be equal to 0 at the cut-off using the pair_modify command. The unit system is real, in which energy is in kcal/mol, distance in Ångströms, or time in femtosecond, has been chosen for practical reasons: the WHAM algorithm used in the second part of the tutorial automatically assumes the energy to be in kcal/mol.
@@
The variables 0, , and , defined in the previous subsection, are used here to create the repulsive potential, restricting the atoms from exploring the center of the box:
@@
Fig. 41 shows the potential and force along the -axis. With , 0 is of the same order of magnitude as the thermal energy , where is the Boltzmann constant and is the temperature used in this simulation. Under these conditions, particles are expected to frequently overcome the energy barrier due to thermal agitation.
@@
Here, the chunk/atom command discretizes the simulation domain into spatial bins of size 2 Å along the direction, and the fix ave/chunk command outputs the number density of atoms within each bin to the file free-sampling.dat. The step count is reset to 0 using reset_timestep to synchronize it with the output times of fix ave/chunk. Run the simulation using LAMMPS.
@@
Increasing the value of 0 reduces the average number of atoms in the central region, making it difficult to achieve a high-resolution free energy profile within reasonable simulation times. For example, running the same simulation with , corresponding to , results in no atoms exploring the central part of the simulation box during the simulation. In such a case, employing an enhanced sampling method is recommended, as done in the next section.
@@ 3.7.2 Method 2: Umbrella sampling
The first difference from the previous case is the larger value for the repulsive potential parameter 0, which makes the central area of the system very unlikely to be visited by free particles. The second difference is the introduction of the variable , which will be used for the biasing potential.
@@ 3.8.1 Creating the system
The class2 pair_styles compute a 6/9 Lennard-Jones potential [70]. The class2 bond, angle, dihedral, and improper styles are used as well, see the documentation for a description of the respective potential form they, each, prescribe. The tail yes option adds long-range van der Waals tail corrections to the energy and pressure. The mix sixthpower imposes the following mixing rule for the calculation of the cross coefficients:
@@
These commands were covered in earlier tutorials and should already be familiar.
@@ 3.8.3 Simulating the reaction
Here, the thermo_style custom command is used to print the cumulative reaction counts which are calculated by fix rxn and thus can be extracted from it. Run the simulation using LAMMPS. As the simulation progresses, polymer chains are observed forming (Fig. 49). During this reaction process, the temperature of the system remains well-controlled (Fig. 50 a), while the number of reactions, r, increases with time (Fig. 50 b).